how do you find the answer for the equations y=2x-2 and y=x+2?
for substitution

Answer by sdmmadam@yahoo.com(530)

how do you find the answer for the equations y=2x-2 and y=x+2?
for substitution
y=2x-2 ----(1)
y=x+2 ----(2)
(putting (1) in (2) that is substituting for y)
2x-2 = x+2
2x-x = 2+2 (grouping like terms, changing sign while changing side)
x = 4
x=4 in (2) implies
y = x+2 = 4+2 = 6
Answer: x=4 and y=6
Verification: x=4 and y=6in (1)gives (6) = 2X(4)-2 which is correct